The following information is available for Marin Inc. for three recent fiscal years.

2017                                       2016                                       2015

Inventory                            $546,328                              $571,700                              $326,238

Net sales                           1,934,372                             1,695,980                             1,327,594

Cost of goods sold            1,554,675                             1,312,366                             964,008

Calculate the inventory turnover, days in inventory, and gross profit rate for 2017 and 2016. (Round inventory turnover to 1 decimal place, e.g. 5.2, days in inventory to 0 decimal places, e.g. 125 and gross profit rate to 1 decimal place, e.g. 5.2%.)
Answer #1

For 2016

Gross profit = Net Sales - Cost of goods sold

= 1,695,980-1,312,366

= $383,614

Gross profit rate = Gross profit / Net sales

= 383,614/1,695,980

= 22.6%

Average inventory = ( Beginning inventory+Ending inventory)/2

= (326,238+571,700)/2

= $448,969

Inventory turnover = Cost of goods sold/Average inventory

= 1,312,366/448,969

= 2.9

Days in inventory = 365/Inventory turnover

= 365/2.9

= 126 days

For 2017

Gross profit = Net Sales - Cost of goods sold

= 1,934,372-1,554,675   

= $379,697

Gross profit rate = Gross profit / Net sales

= 379,697/1,934,372  

= 19.6%

Average inventory = ( Beginning inventory+Ending inventory)/2

= (571,700+546,328)/2

= $559,014

Inventory turnover = Cost of goods sold/Average inventory

= 1,554,675 /559,014

= 2.8

Days in inventory = 365/Inventory turnover

= 365/2.8

= 130 days

Kindly comment if you need further assistance.

Thanks‼!